OpenStudy (anonymous):

solve 5x + 6 = 2x – 3

OpenStudy (anonymous):

First of all we have : \(5x + 6 = 2x -3\) Transpose (-3) to Left hand side, do you know how ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

see : I have 5x +6 = 2x -3, i want to remove -3 from right hand side so I would add 3 both sides: \(5x + 6 + 3 = 2x - 3+3\) Can you tell me what is -3+3?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

No! See... -3 + 3 = 3 -3 what is 3-3 ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

yes so -3+3 is also __ zero

OpenStudy (anonymous):

Now we have : \(5x + 6 +3 = 2x \) \(5x + 9 = 2x\) Now we will remove 5x from Left hand side, by subtracting 5x both sides.

OpenStudy (anonymous):

\(5x + 9 -5x = 2x - 5x\) What is 2x - 5x ?

OpenStudy (anonymous):

i don't get it math gives me a headache and throws me off

OpenStudy (anonymous):

I have 5x + 9 = 2x Now I would subtract 5x both sides \(\color{blue}{(5x+9)} - \color{red}{5x} = \color{blue}{2x} - \color{red}{5x}\) getting it?

OpenStudy (anonymous):

:/ no

OpenStudy (anonymous):

You have two apples in one box and 3 apples in other box. What if you want to put those 2 apples to the second box? What would u do?

OpenStudy (anonymous):

im completly lost

OpenStudy (anonymous):

Prefer this : http://www.purplemath.com/modules/solvelin.htm

OpenStudy (anonymous):

what do you mean solve this? finding x? or simplyfying this equation?

OpenStudy (anonymous):

5x + 9 - 5x = 2x - 5x (combine like terms) 9 = - 3x (now divide by -3) 9/-3 = -3x/-3 (combine like terms) -3 = x check.. 5x + 9 - 5x = 2x - 5x 5(-3) +9 - 5(-3) = 2(-3) - 5(-3) -15 + 9 + 15 = - 6 + 15 9 = 9 (correct) x = - 3 :)

OpenStudy (anonymous):